### v1

@@ 2 Preliminaries
In this section, we recall some definitions and some related observations. A sequence of distinct vertices forms a path $P$ if any two consecutive vertices are adjacent. Whenever we fix a path $P$ of $G$, we shall refer to the subgraph formed by the edges between the consecutive vertices of $P$. The length of a path $P$, denoted by $|P|$, is the number of its vertices minus one. A path is induced if there are no graph edges joining non-consecutive vertices. A path is isometric if it is a shortest path between its end-vertices. For two vertices $u,v$ of a graph $G$, $\mathsf{d}\left(u,v\right)$ denotes the length of an isometric path between $u$ and $v$. An isometric path $P$ is $v$-rooted if one of the end-vertices of $P$ is $v$.
For a path $P$ of a graph $G$ between two vertices $u$ and $v$, the vertices $V(P)\setminus\{u,v\}$ are internal vertices of $P$. A path between two vertices $u$ and $v$ is called a $(u,v)$-path. Similarly, we have the notions of isometric $(u,v)$-path and induced $(u,v)$-path. For a vertex $r$ of $G$ and a set $S$ of vertices of $G$, the distance of $S$ from $r$, denoted as $\mathsf{d}\left(r,S\right)$, is the minimum of the distance between any vertex of $S$ and $r$. For a subgraph $H$ of $G$, the distance of $H$ w.r.t. $r$ is $\mathsf{d}\left(r,V(H)\right)$. Formally, we have $\mathsf{d}\left(r,S\right)=\min\{\mathsf{d}\left(r,v\right)\colon v\in S\}$ and $\mathsf{d}\left(r,H\right)=\mathsf{d}\left(r,V(H)\right)$. For two sets $S$ and $T$, $\mathsf{d}\left(S,T\right)=\min\{\mathsf{d}\left(r,T\right)\colon r\in S\}$.
@@ 2.1 Asymptotic minors
The subgraphs in $(B_{v}\colon v\in E(H))$ are the “branch sets” of the model. The subgraphs in $(P_{e}\colon e\in E(H))$ are the “branch paths” of the model. If $G$ contains a $K$-fat minor model of a graph $H$, then we say that $H$ is a $K$-fat minor of $G$.
@@
For an integer $t\leq 1$, if a graph class $\cal G$ is $K_{2,t}$-asymptotic minor-free, then $\cal G$ is $U_{t}$-asymptotic minor-free.
@@
Assume for contradiction, that $\cal G$ is not $U_{t}$-asymptotic minor-free. This implies that for every integer $K$, there exists a graph $G\in\cal G$ containing a $K$-fat minor model $\mathcal{M}_{t}$ of $U_{t}$. Let $\mathcal{M}_{t}=\left(B_{w}\colon w\in V(U_{t})\right)\cup\left(P_{e}\colon e\in E(U_{t})\right)$. We shall construct a $\frac{K}{2}$-fat minor model of $K_{2,t}$. Let $u$ denote the universal vertex of $U_{t}$ and $B_{u}$ denote the corresponding branch set in $\mathcal{M}$. For each edge $uv\in E(U_{t})$, let $Q_{uv}\subset V(P_{uv})$ denote an induced path of length at least $K$ such that $Q_{uv}$ meets $B_{u}$ or $B_{v}$ only at its end-vertices. By definition, such a path always exists. For each edge $uv\in E(U_{t})$, let $x_{v}$ denote the vertex of $Q_{uv}$ which is at distance $\frac{K}{2}$ from both of end-vertices of $Q_{uv}$. Now we define the branch sets and branch paths of a $\frac{K}{2}$-fat minor model of a $K_{2,t}$. Define $B^{\prime}_{v}=\{x_{v}\}$. Also define $P^{1}_{v},P^{2}_{v}$ be the two paths obtained by deleting $x_{v}$ from $Q_{uv}$. Define $B^{\prime}_{0}=B_{u}$. Finally let $B^{\prime}_{1}$ be the graph induced by the vertices $\{V(B_{w})\colon w\neq u\in V(U_{t})\}\cup\{V(P_{e})\colon e=ww^{\prime},u\notin\{w,w^{\prime}\}\}$. It is easy to check that $\{B^{\prime}_{0},B^{\prime}_{1}\}\cup\{B^{\prime}_{v}\colon uv\in E(U_{t})\}\cup\{P^{1}_{v},P^{2}_{v}\colon uv\in E(U_{t})\}$ is a $\frac{K}{2}$-minor model of $K_{2,t}$. ∎
@@ 2.2 Isometric path complexity
For a graph $G$ and a vertex $r\in V(G)$, consider the following operations on $G$. First, remove all edges $xy$ from $G$ such that $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)$. Let $G^{\prime}_{r}$ be the resulting graph. Then, for each edge $e=xy\in E(G^{\prime}_{r})$ with $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)-1$, orient $e$ from $y$ to $x$. Let $\overrightarrow{G_{r}}$ be the directed acyclic graph formed after applying the above operation on $G^{\prime}$. Note that this digraph can easily be computed in linear time using a Breadth-First Search (BFS) traversal with starting vertex $r$.
The following definition in[10, 11] is inspired by the terminology of posets (as the graph $\overrightarrow{G_{r}}$ can be seen as the Hasse diagram of a poset). For a graph $G$ and a vertex $r\in V(G)$, two vertices $x,y\in V(G)$ are antichain vertices if there are no directed paths from $x$ to $y$ or from $y$ to $x$ in $\overrightarrow{G_{r}}$. A set $X$ of vertices of $G$ is an antichain set if any two vertices in $X$ are antichain vertices. A dual notion of isometric path complexity was defined in [10].
Let $r$ be a vertex of a graph $G$. For a subgraph $H$, $A_{r}\left(H\right)$ shall denote the maximum antichain set of $H$ in $\overrightarrow{G_{r}}$. Let $ipacc\left(\overrightarrow{G_{r}}\right)=\max\left\{|A_{r}\left(P\right)|\colon~{}P~{}\text{is an isometric path}\right\}.$ Define $ipacc\left(G\right)=\min\left\{ipacc\left(\overrightarrow{G_{r}}\right)\colon r\in V(G)\right\}.$
For a graph $G$ and a vertex $r$ of $G$, let $ipco\left(\overrightarrow{G_{r}}\right)$ denote the minimum integer $k$ such that any isometric path $P$ of $G$ can be covered by $k$ $r$-rooted isometric paths. Using Dilworth’s Theorem, one can prove the following important lemma and proposition.
@@
For any graph $G$ and vertex $r$, $ipco\left(\overrightarrow{G_{r}}\right)=ipacc\left(\overrightarrow{G_{r}}\right)$. Therefore, $ipco\left(G\right)=ipacc\left(G\right)$.
@@
Let $G$ be a graph and $r$, an arbitrary vertex of $G$. Consider the directed acyclic graph $\overrightarrow{G_{r}}$, and let $P$ be an isometric path between two vertices $x$ and $y$ in $G$. Then $|P|\geq|\mathsf{d}\left(r,x\right)-\mathsf{d}\left(r,y\right)|+|A_{r}\left(P\right)|-1$.
@@
Orient the edges of $P$ from $y$ to $x$ in $G$. First, observe that $P$ must contain a set $E_{1}$ of oriented edges such that $|E_{1}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|$ and for any $\overrightarrow{ab}\in E_{1}$, $\mathsf{d}\left(r,a\right)=\mathsf{d}\left(r,b\right)+1$. Let the vertices of the largest antichain set of $P$ in $\overrightarrow{G_{r}}$, i.e., $A_{r}\left(P\right)$, be ordered as $a_{1},a_{2},\ldots,a_{t}$ according to their occurrence while traversing $P$ from $y$ to $x$. For $i\in[2,t]$, let $P_{i}$ be the subpath of $P$ between $a_{i-1}$ and $a_{i}$. Observe that for any $i\in[2,t]$, since $a_{i}$ and $a_{i-1}$ are antichain vertices, there must exist an oriented edge $\overrightarrow{b_{i}c_{i}}\in E(P_{i})$ such that either $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$ or $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)-1$. Let $E_{2}=\{b_{i}c_{i}\}_{i\in[2,t]}$. Observe that $E_{1}\cap E_{2}=\emptyset$ and therefore $|P|\geq|E_{1}|+|E_{2}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|+|A_{r}\left(P\right)|-1$. ∎
@@ 3 Asymptotic minors and graphs with restricted holes
In Section 3.1 we show that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity and prove Theorem 1. In Section 3.2 we show that monoholed graphs are $U_{4}$-asymptotic minor-free and prove Theorem 2. Finally, in Section 3.3 we show that even-hole free graphs have unbounded isometric path complexity and prove Theorem 3.
@@ 3.1 Asymptotic minors
In this section, we prove Theorem 1. Let $\mathcal{G}$ be a graph class that does not contain $U_{t}$ as an asymptotic minor, for some $t\geq 1$. Then, by the definition, there exists an integer $K$ and a graph $G\in\mathcal{G}$ such that $G$ does not contain a $K$-fat minor model of $U_{t}$. We can assume $K\geq 5$. The following lemma and Proposition 2 prove Theorem 1.
Lemma 2.
For any two integers $t\geq 2,K\geq 5$, if $G$ does not contain a $K$-fat minor model of $U_{t}$, then $ipacc\left(G\right)\leq(10K+9)t-1$.
We devote the remainder of this section to proving Lemma 2. For integers $t\geq 2,K\geq 5$, let $G$ be a graph that does not contain a $K$-fat minor model of $U_{t}$, such that $ipacc\left(G\right)\geq(10K+9)t$. First, we introduce some notation that will be used throughout the proof.
Let $r$ be any vertex of $G$. Note that $ipacc\left(\overrightarrow{G_{r}}\right)\geq(10K+9)t$. Hence, there exists an isometric path $P$ with $|A_{r}\left(P\right)|\geq(10K+9)t$. Assume that the two end-vertices of $P$ are $u$ and $v$, respectively. Notice that $r\notin\left\{u,v\right\}$. Let $A=A_{r}\left(P\right)$, $k=|A|$ and the elements of $A$ are ordered as $p_{1},p_{2},\ldots,p_{k}$ as they are encountered while traversing $P$ from $u$ to $v$. Let $a_{1}=p_{1}$, and for $i\in[2,t]$ let $n(i)=(10K+9)(i-1)$, $a_{i}=p_{n(i)}$. For convenience, we redefine $A=\left\{a_{i}\colon i\in[t]\right\}$. Recall that this set contains antichain vertices.
@@
Figure 3: Illustration of the proof of Lemma 2. The shaded triangular shapes indicate the branch sets $Z_{1},Z_{2},\ldots,Z_{t}$, and subpaths of $P,Q_{1},Q_{2},\ldots,Q_{t}$ are used as branch paths to construct a $K$-fat minor model of $U_{t}$.
Input : A root vertex $r$, an isometric path $P$ whose end-vertices are distinct from $r$, a vertex $w\in V(P)$, and an end-vertex $u$ of $P$.
Output : An induced path.
@@
2Let $Q$ be any $(r,w)$-isometric path in $G$;
3Let $Q^{\prime}$ be the maximal $(r,w^{\prime})$-subpath of $Q$ such that no internal vertex of $Q^{\prime}$ is adjacent to any vertex of $P$.
4if $w^{\prime}\in V(P)$ then
@@
7       Let $w^{\prime\prime}\in V(P)\cap N(w^{\prime})$ be the vertex such that no internal vertex of the $(u,w^{\prime\prime})$-subpath of $P$ is a neighbor of $w^{\prime}$.
8      Let $Q^{\prime\prime}$ be the path obtained by appending $w^{\prime\prime}$ to $Q^{\prime}$.
@@
Procedure 1 A procedure $\textsc{Connect}\left(r,P,w,u\right)$ to build pillars.
Finding special induced paths: Now we shall describe a general procedure to find special induced paths which will allow us to create branch paths and branch sets of a $K$-fat minor model of $U_{t}$ (and hence arrive to a contradiction). Consider Procedure 1. Let $Q_{i}$ denote the output of $\textsc{Connect}\left(r,P,a_{i},u\right)$ where $a_{i}\in A$. For each $i\in[t]$, observe that one end-vertex of $Q_{i}$ lies on $P$. Let us denote this end-vertex as $b_{i}$ and its neighbor in $Q_{i}$ as $b^{\prime}_{i}$. We prove the following claims that establish the special properties of these paths.
Claim 5.1.
@@
To prove (a) fix any $i\in[t]$. The path $Q_{i}$ is the path returned by $\textsc{Connect}\left(r,P,a_{i},u\right)$. If $\textsc{Connect}\left(r,P,a_{i},u\right)$ returned from Step-$4$ of Procedure 1, then $Q_{i}$ is an $(r,w_{i})$-isometric path and therefore is an induced path. Otherwise, $\textsc{Connect}\left(r,P,a_{i},u\right)$ must have returned from Step-$8$ of Procedure 1. In this case, $Q^{\prime}_{i}$ is constructed by taking a subpath of a $(r,a_{i})$-isometric path and appending to it one more vertex from the path $P$. By construction (see Step-$1$ and Step-$6$), $Q^{\prime}_{i}$ remains an induced path. Both $(b)$ and $(c)$ follow from the construction. ∎
@@
Claim 5.2.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$. Then $|A_{r}\left(X_{i}\right)|\leq 3$.
@@
Assume for contradiction that there exists an integer $i\in[t]$ such that $|A_{r}\left(X_{i}\right)|\geq 4$. Hence due to Proposition 2, we have that $\mathsf{d}\left(b_{i},a_{i}\right)\geq 3+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$. Let $R$ denote the $(r,b^{\prime}_{i})$-subpath of $Q_{i}$. Due to 5.1(c) we know that there is an $(r,a_{i})$-isometric path $S$ such that $V(R)\subseteq V(S)$. Hence $b^{\prime}_{i}\in V(S)$ and let $S^{\prime}$ denote the $(b^{\prime}_{i},a_{i})$-subpath of $S$. Observe that appending $b_{i}$ to $S^{\prime}$ gives a path from $b_{i}$ to $a_{i}$. Hence, $\mathsf{d}\left(b_{i},a_{i}\right)\leq 2+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$. But this is a contradiction. ∎
Claim 5.3.
For $\{i,j\}\subseteq[t]$, let $R_{ij}$ denote the $(b_{i},b_{j})$-subpath of $P$. Then $|A_{r}\left(R_{ij}\right)|\geq(10K+3)$.

### v2

@@ 2 Preliminaries
In this section, we recall some definitions and some related observations. A sequence of distinct vertices form a path $P$, if any two consecutive vertices are adjacent. Whenever we fix a path $P$ in $G$, we shall refer to the subgraph formed by the edges between the consecutive vertices of $P$. The length of a path $P$ is the number of edges in $P$. A path is induced if there are no graph edges joining non-consecutive vertices. A path is isometric if it is a shortest path between its end-vertices. For two vertices $u,v$ of a graph $G$, $\mathsf{d}\left(u,v\right)$ denotes the length of an isometric path between $u$ and $v$. An isometric path $P$ is $v$-rooted if one of the end-vertices of $P$ is $v$.
For a path $P$ of a graph $G$ between two vertices $u$ and $v$, the vertices $V(P)\setminus\{u,v\}$ are internal vertices of $P$. A path between two vertices $u$ and $v$ is called a $(u,v)$-path. Similarly, we have the notions of isometric $(u,v)$-path and induced $(u,v)$-path. For a vertex $r$ of $G$ and a set $S$ of vertices of $G$, the distance of $S$ from $r$, denoted as $\mathsf{d}\left(r,S\right)$, is the minimum of the distances between any vertex of $S$ from $r$, that is, $\mathsf{d}\left(r,S\right)=\min\{\mathsf{d}\left(r,v\right)\colon v\in S\}$. For a subgraph $H$ of $G$, the distance of $H$ from $r$ is $\mathsf{d}\left(r,V(H)\right)$. For two sets $S$ and $T$, $\mathsf{d}\left(S,T\right)=\min\{\mathsf{d}\left(r,T\right)\colon r\in S\}$.
@@ 2.1 Asymptotic minors
The subgraphs in $(B_{v}\colon v\in E(H))$ are the “branch sets” of the model. The subgraphs in $(P_{e}\colon e\in E(H))$ are the “branch paths” of the model. If $G$ contains a $K$-fat minor model of a graph $H$, then we say that $H$ is a $K$-fat minor of $G$. Note that, for $e\in E(H)$, we can assume that the connected subgraphs $P_{e}$ are indeed paths.
@@
$B_{u}$$B^{\prime}_{0}$$X_{v}$$B^{\prime}_{1}$
Figure 3: Illustration of the proof of Proposition˜11. Dashed ellipses denote the branch sets of $\mathcal{M}_{t}$. The solid lines indicate the branch paths of $\mathcal{M}_{t}$. The dashed rectangular boxes indicate the branch sets of the $\left\lfloor\frac{K}{4}\right\rfloor$-fat minor model of $K_{2,t}$.
@@
For an integer $t\geq 1$, if a graph class $\cal G$ is $K_{2,t}$-asymptotic minor-free, then $\cal G$ is $U_{t}$-asymptotic minor-free.
@@
See Figure˜3 for illustration. Assume for contradiction, that $\cal G$ is not $U_{t}$-asymptotic minor-free. This implies that for every integer $K\geq 12$, there exists a graph $G\in\cal G$ containing a $K$-fat minor model $\mathcal{M}_{t}$ of $U_{t}$. Let $\mathcal{M}_{t}=\left(B_{w}\colon w\in V(U_{t})\right)\cup\left(P_{e}\colon e\in E(U_{t})\right)$. We shall construct a $\left\lfloor\frac{K}{4}\right\rfloor$-fat minor model of $K_{2,t}$. Let $u$ denote the universal vertex of $U_{t}$ and $B_{u}$ denote the corresponding branch set in $\mathcal{M}_{t}$. For each edge $uv\in E(U_{t})$, let $Q_{uv}\subset V(P_{uv})$ denote an induced path of length at least $K$ such that $Q_{uv}$ meets $B_{u}$ or $B_{v}$ only at its end-vertices. By definition, such a path always exists. For each edge $uv\in E(U_{t})$, let $X_{v}$ denote the subpath of $Q_{uv}$ whose length is exactly $\left\lfloor\frac{K}{4}\right\rfloor+3$ and is at distance at least $\left\lfloor\frac{K}{4}\right\rfloor$ from both of the end-vertices of $Q_{uv}$. Define $P^{1}_{v},P^{2}_{v}$ to be the two paths obtained by deleting the vertices of $X_{v}$ from $Q_{uv}$. Define $B^{\prime}_{0}=B_{u}$. Finally, let $B^{\prime}_{1}$ be the graph induced by the vertices
$\left(\bigcup\limits_{w\in V(U_{t})\setminus\{u\}}V(B_{w})\right)\cup\left(\displaystyle\bigcup\limits_{e\in E(U_{t}-\{u\})}V(P_{e})\right)$
It is easy to check that $\{B^{\prime}_{0},B^{\prime}_{1}\}\cup\{X_{v}\colon uv\in E(U_{t})\}\cup\{P^{1}_{v},P^{2}_{v}\colon uv\in E(U_{t})\}$ is a $\left\lfloor\frac{K}{4}\right\rfloor$-fat minor model of $K_{2,t}$. ∎
@@ 2.2 Isometric path complexity
Let $G$ be a graph, and $r\in V(G)$ be a vertex. In this section, $\mathsf{d}\left(.,.\right)$ refers to the metric in $G$. Let $G_{r}$ be the graph obtained by removing all edges $xy$ from $G$ such that $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)$. Then, for each edge $e=xy\in E(G_{r})$ with $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,y\right)-1$, orient $e$ from $y$ to $x$. Let $\overrightarrow{G_{r}}$ be the directed acyclic graph formed after applying the above operation on $G^{\prime}$. Note that this digraph can easily be computed in linear time using a Breadth-First Search (BFS) traversal with starting vertex $r$.
The following definition in [17, 14] is inspired by the terminology of posets (as the graph $\overrightarrow{G_{r}}$ can be seen as the Hasse diagram of a poset). For a graph $G$ and a vertex $r\in V(G)$, two vertices $x,y\in V(G)$ are antichain vertices if there are no directed paths from $x$ to $y$ or from $y$ to $x$ in $\overrightarrow{G_{r}}$. A set $X$ of vertices of $G$ is an antichain set if any two vertices in $X$ are antichain vertices. For a subgraph $H$ of $G$, $A_{r}\left(H\right)$ shall denote a maximum antichain set of $H$ in $\overrightarrow{G_{r}}$. Let $ipacc\left(\overrightarrow{G_{r}}\right)=\max\left\{|A_{r}\left(P\right)|\colon~P~\text{is an isometric path in }G\right\}.$ Define $ipacc\left(G\right)=\min\left\{ipacc\left(\overrightarrow{G_{r}}\right)\colon r\in V(G)\right\}.$ Let $ipco\left(G,r\right)$ denote the minimum integer $k$ such that any isometric path $P$ of $G$ can be covered by $k$ many $r$-rooted isometric paths in $G$. Recall that, $ipco\left(G\right)=\min\{ipco\left(G,r\right)\colon r\in V(G)\}$ and $sipco\left(G\right)=\max\{ipco\left(G,r\right)\colon r\in V(G)\}$. Using Dilworth’s Theorem, one can prove the following important lemma and proposition.
@@
For any graph $G$ and vertex $r$, $ipco\left(G,r\right)=ipacc\left(\overrightarrow{G_{r}}\right)$. Therefore, $ipco\left(G\right)=ipacc\left(G\right)$.
@@
Let $G$ be a graph and $r$, an arbitrary vertex of $G$. Consider the directed acyclic graph $\overrightarrow{G_{r}}$, and let $P$ be an isometric path between two vertices $x$ and $y$ in $G$. Then $|E(P)|\geq|\mathsf{d}\left(r,x\right)-\mathsf{d}\left(r,y\right)|+|A_{r}\left(P\right)|-1$.
@@
Orient the edges of $P$ from $y$ to $x$ in $G$. First, observe that $P$ must contain a set $E_{1}$ of oriented edges such that $|E_{1}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|$ and for any $\overrightarrow{ab}\in E_{1}$, $\mathsf{d}\left(r,a\right)=\mathsf{d}\left(r,b\right)+1$. Let the vertices of a largest antichain set of $P$ in $\overrightarrow{G_{r}}$, i.e., $A_{r}\left(P\right)$, be ordered as $a_{1},a_{2},\ldots,a_{t}$ according to their occurrence while traversing $P$ from $y$ to $x$. For $i\in[2,t]$, let $P_{i}$ be the subpath of $P$ between $a_{i-1}$ and $a_{i}$. Observe that for any $i\in[2,t]$, since $a_{i}$ and $a_{i-1}$ are antichain vertices, there must exist an oriented edge $\overrightarrow{b_{i}c_{i}}\in E(P_{i})$ such that either $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)$ or $\mathsf{d}\left(r,b_{i}\right)=\mathsf{d}\left(r,c_{i}\right)-1$. Let $E_{2}=\{b_{i}c_{i}\}_{i\in[2,t]}$. Observe that $E_{1}\cap E_{2}=\emptyset$ and therefore $|E(P)|\geq|E_{1}|+|E_{2}|=|\mathsf{d}\left(r,y\right)-\mathsf{d}\left(r,x\right)|+|A_{r}\left(P\right)|-1$. ∎
@@ 3 Asymptotic minors and graphs with restricted holes
In Section˜3.1, we show that $U_{t}$-asymptotic minor-free graphs have bounded strong isometric path complexity and prove Theorem˜2. In Section˜3.2, we show that monoholed graphs are $U_{4}$-asymptotic minor-free and prove Theorem˜6. Finally, in Section˜3.3, we show that even-hole free graphs have unbounded isometric path complexity and prove Theorem˜7.
@@ 3.1 Asymptotic minors
In this section, we prove Theorem˜2. Let $\mathcal{G}$ be a graph class that does not contain $U_{t}$ as an asymptotic minor, for some $t\geq 1$. Then, by the definition, there exists an integer $K$ and a graph $G\in\mathcal{G}$ such that $G$ does not contain a $K$-fat minor model of $U_{t}$. We can assume $K\geq 5$. The following and Lemma˜12 prove Theorem˜2.
Theorem 14.
For any two integers $t\geq 2,K\geq 5$, if $G$ does not contain a $K$-fat minor model of $U_{t}$, then for any vertex $r\in V(G)$, $ipacc\left(\overrightarrow{G_{r}}\right)\leq({12K}+{5})t-1$.
We devote the remainder of this section to proving Theorem˜14. For integers $t\geq 2,K\geq 5$, assume for contradiction that there exists a graph $G$ that does not contain a $K$-fat minor model of $U_{t}$, and there exists a vertex $r\in V(G)$ such that $ipacc\left(\overrightarrow{G_{r}}\right)\geq({12K}+{5})t$.
Observe that there exists an isometric path $P$ with $|A_{r}\left(P\right)|\geq({12K}+{5})t$. Assume that the two end-vertices of $P$ are $u$ and $v$, respectively. Notice that $r\notin\left\{u,v\right\}$ (otherwise we would have $A_{r}\left(P\right)=\emptyset$). Let $k=|A_{r}\left(P\right)|$ and the elements of $A_{r}\left(P\right)$ are ordered as $p_{1},p_{2},\ldots,p_{k}$ as they are encountered while traversing $P$ from $u$ to $v$. Let $a_{1}=p_{1}$, and for $i\in[2,t]$ let n(i)=(12K+5)(i1)n(i)=({12K}+{5})(i-1), $a_{i}=p_{n(i)}$. Let $A=\left\{a_{i}\colon i\in[t]\right\}$. Recall that this set forms an antichain in $\overrightarrow{G_{r}}$, and its definition implies that any two of its vertices are far apart from each other.
@@
Figure 4: Illustration of the notations used to denote the paths returned by Procedure 1 for the horizontal $(u,v)$-path $P$ (indicated by the solid black line). The path $Q_{1}$ (indicated with dash-dots) is the vertical $(r,b^{\prime}_{1})$-path with $b_{1}$ appended to it. The path $Q_{2}$ (indicated with dots) is the downward $(r,a_{2})$-path.
Input : A root vertex $r$, an isometric path $P$ whose end-vertices are distinct from $r$, a vertex $a\in V(P)$, and an end-vertex $u$ of $P$.
Output : A path.
@@
2Let $Q$ be any $(r,a)$-isometric path in $G$;
3Let $Q^{\prime}$ be the maximal $(r,a^{\prime})$-subpath of $Q$ such that no internal vertex of $Q^{\prime}$ is adjacent to any vertex of $P$.
4if $a^{\prime}\in V(P)$ then
@@
7   Let $a^{\prime\prime}\in V(P)\cap N(a^{\prime})$ be the vertex such that no internal vertex of the $(u,a^{\prime\prime})$-subpath of $P$ is a neighbor of $a^{\prime}$.
8  Let $Q^{\prime\prime}$ be the path obtained by appending $a^{\prime\prime}$ to $Q^{\prime}$.
@@
Procedure 1 A procedure $\textsc{Connect}\left(r,P,a,u\right)$ to build pillars.
Now, using the vertices of $A$ and Procedure 1, we construct special induced paths that will aid us to construct the $K$-fat minor model of $U_{t}$. First, we illustrate Procedure 1 with an example.
Illustration of Procedure 1. For illustration purposes, consider the graph in Figure˜4. Let $P$ be the $(u,v)$-path shown in the figure, and consider $\textsc{Connect}\left(r,P,a_{1},u\right)$. In this case, the vertex $a$ (used in the description of Proposition 1) is $a_{1}$, Procedure 1 returns from Line 8, and therefore $Q_{1}=r\ldots b^{\prime}_{1}b_{1}$. Now consider $\textsc{Connect}\left(r,P,a_{1},u\right)$. In this case, Procedure 1 returns from Line 4, $b_{2}=a_{2}$, and $Q_{2}$ is the downward $(r,a_{2})$-path.
Finding special induced paths. Now, we shall describe a general procedure to find special induced paths which will allow us to create branch paths and branch sets of a $K$-fat minor model of $U_{t}$ (and hence arrive to a contradiction). Consider Procedure 1. Let $Q_{i}$ denote the output of $\textsc{Connect}\left(r,P,a_{i},u\right)$ where $a_{i}\in A$. For each $i\in[t]$, observe that one end-vertex of $Q_{i}$ lies on $P$. Let us denote this end-vertex as $b_{i}$ and its neighbor in $Q_{i}$ as $b^{\prime}_{i}$. The construction of Procedure 1 implies the following claim establishing some special properties of these paths.
Claim 3.1.A.
@@
To prove (a) fix any $i\in[t]$. The path $Q_{i}$ is the path returned by $\textsc{Connect}\left(r,P,a_{i},u\right)$. If $\textsc{Connect}\left(r,P,a_{i},u\right)$ returned from Step-$4$ of Procedure 1, then $Q_{i}$ is an $(r,w_{i})$-isometric path and therefore is an induced path. Otherwise, $\textsc{Connect}\left(r,P,a_{i},u\right)$ must have returned from Step-$8$ of Procedure 1. In this case, $Q^{\prime}_{i}$ is constructed by taking a subpath of an $(r,a_{i})$-isometric path and appending to it one more vertex from the path $P$. By construction (see Step-$1$ and Step-$6$), $Q^{\prime}_{i}$ remains an induced path. $(b)$ follows easily from the construction. To observe $(c)$, we consider two cases. If Procedure 1 returned from Line 4, then Line 1 and 2 imply $b_{i}=a_{i}$, and the path returned by $\textsc{Connect}\left(r,P,a_{i},u\right)$ is itself an $(r,a_{i})$-isometric path and thus (c) follows. Otherwise, $b_{i}\not=a_{i}$ but lies on $P$, and due to Line 6 and 7, $Q_{i}-b_{i}$ is a subpath of the $(r,a_{i})$-isometric path considered in Line 1. Now (c) follows. ∎
@@
[…]
(a) Illustration of the proof of ˜3.1.B. The rectilinear curve (made of solid and dashed lines) between $u$ and $v$ represents $P$. The (vertical) dotted line represents $R$. The (vertical) dash-dotted line represents $S^{\prime}$ .
[…]
(b) Illustration of the proof of ˜3.1.C. The dotted line between $r,b_{j}$ indicates $Q_{j}$. The dotted line between $r,b_{i}$ indicates $Q_{i}$. The gray line and the rectilinear black curve together indicate $S$. The gray line also indicates $R_{ij}$ in this example.
[…]
(c) Illustration of the proof of ˜3.1.D. The thick gray line indicates $Q^{\prime}_{i}$. The dotted line between $r$ and $a_{j}$ indicates $Q_{j}$. The dotted line between $r$ and $c_{i}$ indicates a subpath of $Q_{i}$. The gray line indicates $Q^{\prime}_{i}$. The dashed line indicates a $(c^{\prime}_{i},d_{j})$ isometric path.
[…]
(d) Illustration of the proof of ˜3.1.E. The dotted line along with the vertex labeled $b_{i}$ indicates $Q_{i}$. The gray line indicates a $(c,d)$-isometric path.
Figure 5: Illustrations of notations used in the proofs for the claims in Theorem˜14.
Claim 3.1.B.
For each $i\in[t]$, let $X_{i}$ denote the $(a_{i},b_{i})$-subpath of $P$. Then $|A\cap V(X_{i})|\leq 3$.
@@
See Figure˜5(a). Let $R$ denote the $(r,b^{\prime}_{i})$-subpath of $Q_{i}$. Due to ˜3.1.A(c), we know that there is an $(r,a_{i})$-isometric path $S$ such that $V(R)\subseteq V(S)$. Hence, $b^{\prime}_{i}\in V(S)$; let $S^{\prime}$ denote the $(b^{\prime}_{i},a_{i})$-subpath of $S$. Observe that appending $b_{i}$ to $S^{\prime}$ gives a path from $b_{i}$ to $a_{i}$. Since $\mathsf{d}\left(b^{\prime}_{i},a_{i}\right)=|\mathsf{d}\left(r,a_{i}\right)-\mathsf{d}\left(r,b^{\prime}_{i}\right)|$ and $|\mathsf{d}\left(r,b^{\prime}_{i}\right)-\mathsf{d}\left(r,b_{i}\right)|\leq 1$, we have, $\mathsf{d}\left(b_{i},a_{i}\right)\leq 2+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$. If $|A\cap V(X_{i})|\geq 4$, then Proposition˜13 implies $\mathsf{d}\left(b_{i},a_{i}\right)\geq 3+|\mathsf{d}\left(r,b_{i}\right)-\mathsf{d}\left(r,a_{i}\right)|$, a contradiction. Hence, $|A\cap V(X_{i})|\leq 3$. ∎
Claim 3.1.C.
For $\{i,j\}\subseteq[t]$, let $R_{ij}$ denote the $(b_{i},b_{j})$-subpath of $P$. Then $|A\cap V(R_{ij})|\geq{12K}+2$.